1- A space heater operates on 120 VAC RMS household current. The RMS current that passes through the space heater is 8 amps. If the current and the voltage are 30 degrees out of phase, how much power does the space heater use on average?

A) 831 Watts                   b) 960 Watts                      c) 148 Watts          d) 120 Watts     e) none of these

2. given that C=2 farads and R= 1000 Ohms, about how long will it take the capacitor in the circuit below to charge once the switch is closed?

A) 11.1 hours                b) 20 seconds              c)1.11 hours          d) 2000 seconds  
Answer #1

. 1.) p = Vamos Irms casa - (1200)*(8A) * cos(30°) P = 831 w 2) T=RC = (1000-2) 2F) T = 2000s